Estimate the wavelength of x-rays emitted by platinum as an electron falls from n = 2 to n = 1. Platinum has Z = 78. n


1
Expert's answer
2021-12-06T09:45:37-0500

By the Rydberg's Equation:


"\frac1λ = RZ^2\bigg(\frac1{n_1^2} - \frac1{n_2^2} \bigg),\\\space\\
\lambda=\frac{(n_1n_2)^2}{RZ^2(n_2^2-n_1^2)}=20\text{ pm}."